The following is the age distribution of 1000 working employees working in an organization

Age(years) 20-25 25-30 30-35 35-40 40-45 45-50 50-55 55-60 60-65
Number of persons 30 160 210 180 145 105 70 60 40

Due to continuous losses, it is desired to bring down the manpower according the following scheme

1. Retrench the first 15% from the lower age group

2. Absorb the next 45% in the other branches.

3. Make 10% from the highest age group permanently retire, if necessary.

a) Calculate the age limits of the persons retained and those transferred to other departments.

b) Find the average age of those retained and those transferred to other branches

c) Calculate the standard deviation of the employee retained and transferred

d) Compute the coefficient of skewness of the employees retained and transferred

Comment on your result
Answer #1
Age(years) 20-25 25-30 30-35 35-40 40-45 45-50 50-55 55-60 60-65
Number of persons 30 160 210 180 145 105 70 60 40
cumulative 30 190 400 580 725 830 900 960 1000

we are retrenching, that is, laying off first 15% from the lower age group

thus 150 of the 1000 employee are cut off starting from the lower age group

so the lowest income group is gone.

and we are left with the following

Age(years) 25-30 30-35 35-40 40-45 45-50 50-55 55-60 60-65
Number of persons 40 210 180 145 105 70 60 40

the next 45% is transferred

thus 450 of the next employees are transfered which looks like this

Age(years) 25-30 30-35 35-40 40-45
Number of persons 40 210 180 20

and we are left with the following

Age(years) 40-45 45-50 50-55 55-60 60-65
Number of persons 125 105 70 60 40

also we are permanently retiring 10% from the highest age group,that is 4 persons, so we are left with

Age(years) 40-45 45-50 50-55 55-60 60-65
Number of persons 125 105 70 60 36

a)the transfered employees are

Age(years) 25-30 30-35 35-40 40-45
Number of persons 40 210 180 20

the retained employees are

Age(years) 40-45 45-50 50-55 55-60 60-65
Number of persons 125 105 70 60 36

b)for finding the average we calculate the midpoint of every class and we get the following

Age(years) 20-25 25-30 30-35 35-40 40-45 45-50 50-55 55-60 60-65
mid point 22.5 27.5 32.5 37.5 42.5 47.5 52.5 57.5 62.5

Thus the average=\sum (mid\ point*frequency\ of\ that\ class)/total\ frequency

therefore average of the transfered employee= 34.5

average of the retained employees=49.68

c)Standard deviation=\sqrt{1/n\sum(m.p-mean)^2*freq} , m.p=mid point of class, n=total frequency

therefore standard deviation of retained employees=6.56

standard deviation of transfered employees=3.56

d)skewness=mean-mode/standard\ deviation

skewness of retained employee=(49.68-42.5)/6.56 = 1.09

skewness of transfered employee=(34.5-32.5)/3.56= 0.56

we can see that both the data set are positively skewed with retained employee being more positively skewed than transfered employee and transfered employee is quite close to symmetry